# Minimal plus-one generated curves and Briançon-type polynomials

### v1

@@ Abstract.
We define the minimal plus-one generated curves and prove a result explaining why they are the closest relatives of the free curves, after the nearly free curves. Then we look at the projective closures of the general and of the special fibers of some Briançon-type polynomials constructed by E. Artal Bartolo, Pi. Cassou-Noguès and I. Luengo Velasco. They yield new examples of free, nearly free or minimal plus-one generated curves. This leads us to ask the natural question if such a result holds for any Briançon-type polynomials.
@@ 1. Introduction
In this note we first define the minimal plus-one generated curves in Definition 1.1 and give a simple numerical characterization of them in Theorem 1.2. This characterization is similar to the known characterizations of free and nearly free curves, which are recalled in Corollaries 2.2 and 2.3. Then we construct new examples of free, nearly free and minimal plus-one generated curves starting with some Briançon-type polynomials constructed in [3].
Let us recall the definitions of all these objects. Let $S=\mathbb{C}[x,y,z]$ be the polynomial ring in three variables $x,y,z$ with complex coefficients, and let $C:f=0$ be a reduced curve of degree $d\geq 3$ in the complex projective plane $\mathbb{P}^{2}$. We denote by $J_{f}$ the Jacobian ideal of $f$, i.e., the homogeneous ideal in $S$ spanned by the partial derivatives $f_{x},f_{y},f_{z}$ of $f$, and by $M(f)=S/J_{f}$ the corresponding graded quotient ring, called the Jacobian (or Milnor) algebra of $f$. Consider the graded $S$-module of Jacobian syzygies of $f$ or, equivalently, the module of derivations killing $f$, namely
@@
We say that $C:f=0$ is an $m$-syzygy curve if the module $D_{0}(f)$ is minimally generated by $m$ homogeneous syzygies, say $\rho_{1},\rho_{2},\ldots,\rho_{m}$, of degrees $d_{j}=\deg\rho_{j}$ ordered such that
@@
The curve $C$ is free when $m=2$, since then $D_{0}(f)$ is a free module of rank 2, see for instance [22, 23, 24, 25]. A free curve $C$ is also characterized by the condition $d_{1}+d_{2}=d-1$, and hence in this case $r=d_{1}<d/2$, see [24]. A nearly free curve is a 3-syzygy curve $C$ with exponents $(d_{1},d_{2},d_{3})$ such that $d_{1}+d_{2}=d$ and $d_{2}=d_{3}$. To simplify the notation, we say that $(d_{1},d_{2})$ are the exponents for the nearly free curve $C$. Finally, a plus-one generated curve is a 3-syzygy curve $C$ with exponents $(d_{1},d_{2},d_{3})$ such that $d_{1}+d_{2}=d$, see [1, 15]. We define the $\delta$-level of a plus-one generated curve by the formula
@@
These curves can be regarded as the closest neighbors of the free curves beyond the nearly free curves, as shown by our main result in this note.
Theorem 1.2.
$\tau(C)=\tau(d,r)_{max}-2.$
$N(f)=I_{f}/J_{f}.$
$\nu(C)=\max_{j}\{n(f)_{j}\}$
as in [2]. Note that $C$ is free if and only if $N(f)=0$, see for instance [24], and hence in this case $\nu(C)=0$, and $C$ is nearly free if and only if $\nu(C)=1$, see [13]. To prove Theorem 1.2, we show in fact that $\tau(C)=\tau(d,r)_{max}-2$ implies $\nu(C)=2$ and then use the classification of such curves given in [15, Theorem 3.11]. This result says that a curve $C$ with $\nu(C)=2$ is either an MPOG-curve, or a 4-syzygy curve with exponents $(r,r,r,r)$ and degree $d=2r-1$, which is in fact a maximal Tjurina curve of type $(d,r)=(2r-1,r)$ as defined in [16], see also Section 2 below.
Let $s=xy+1$, $p=xs+1$ and $u=s^{2}+y$. Using this notation, we define
The polynomial $g$ up-to a change of coordinates $(x,y)\mapsto(-x,-y)$ was considered in an unpublished manuscript by J. Briançon, see [3, 4]. We recall that for a polynomial mapping $h:\mathbb{C}^{2}\to\mathbb{C}$ the set of atypical values is the minimal set $A(h)\subset\mathbb{C}$ such that $h$ induces a locally trivial topological fibration over the set $\mathbb{C}\setminus A(h)\cup C(h)$, where $C(h)$ denotes the set of critical values of $h$. The atypical values for a function $h:\mathbb{C}^{2}\to\mathbb{C}$ come from jumping of the Milnor numbers at infinity, see [5, 18, 20].
The polynomials $g$ and $g^{\prime}$ have the following very interesting properties, see [3].
Theorem 1.3.
  1. (1)
  2. (2)
  3. (3)
    The set of atypical values for $g$ is $A(g)=\{0,b\}$, where $b=-\frac{16}{9}.$
  4. (4)
  5. (5)
When $h:\mathbb{C}^{2}\to\mathbb{C}$ is obtained from a linear form in $R$ by composition with a polynomial automorphism of $\mathbb{C}^{2}$, the freeness properties of the projective closure $D_{t}$ of any fiber $h^{-1}(t)$ are discussed in [11, Theorem 1.5]. It is natural to ask what happens when we replace such a polynomial $h$ by a Briançon-type polynomial.
For the polynomial $g$, we consider three associated plane curves $C_{t}$ in $\mathbb{P}^{2}$, which are the projective closure of the affine fibers $F_{t}=g^{-1}(t)$ for $t=0$, $t=b$ and $t=1$. The first two fibers correspond to the atypical values in $A(g)$, and the third one is a general fiber of $g$. For the convenience of the reader, the defining equations for these curves, obtained by homogeneization starting with the first equation in (1.3), are given in (3.1) and (3.2). Similarly, for the polynomial $g^{\prime}$, we consider three associated plane curves $C_{t}^{\prime}$ in $\mathbb{P}^{2}$, which are the projective closure of the affine fibers $g^{\prime-1}(t)$ for $t=0$, $t=b^{\prime}$ and $t=1$. The corresponding equations are given in (3.3) and (3.4). Here are our results about the freeness properties of these projective plane curves.
Theorem 1.4.
The geometric genera of the curves $C_{t}$ are given by
$g(C_{1})=1\text{ and }g(C_{0})=g(C_{b})=0.$
  1. (1)
  2. (2)
    The curve $C_{b}$ is a free curve with exponents $(4,5)$ and $\tau(C_{b})=61$.
Theorem 1.5.
The geometric genera of the curves $C_{t}^{\prime}$ are given by
  1. (1)
  2. (2)
Remark 1.6.
The claims in Theorem 1.4 and Theorem 1.5 about the genera of the curves $C_{t}$ and $C_{t}^{\prime}$ are already proved in [3]. We state them here for the following reason. Among the six curves considered in Theorem 1.4 and Theorem 1.5, the curve $C_{b}$ is rational cuspidal and the curves $C_{0}^{\prime}$ and $C_{b^{\prime}}^{\prime}$ are rational nearly cuspidal curves, that is all their singularities are irreducible except one which has 2 branches, see Remarks 3.1 and 3.2. Hence the claims above are compatible with fact that a rational cuspidal curve of even degree is either free or nearly free, see [13], and a rational nearly cuspidal curve $C$ of even degree is free, nearly free or an $MPOG$-curve, see [15, 13]. The curve $C_{0}$, which is not cuspidal, shows moreover that the hypothesis of being cuspidal is necessary in general for the first result to hold, even if there are exceptions, as shown by the curves $C_{0}^{\prime}$ and $C_{b^{\prime}}^{\prime}$ which are not cuspidal. The curves $C_{1}$ and $C_{1}^{\prime}$ are both $MPOG$-curves, even though they are not rational.
Since the line at infinity $L_{z}:z=0$ intersects the curves $C_{t}$ only in two points, the general principle explained in [12] tells us that the curves $C_{t}\cup L_{z}$ (resp. $C_{t}^{\prime}\cup L_{z}$) of degree $11$ may also enjoy (sometimes better) freeness properties. See also Step 3 in the proof of Theorem 1.5 in [11] for such a situation. We consider also the curve arrangements $C_{0}\cup C_{b}$ (resp. $C_{0}^{\prime}\cup C^{\prime}_{b^{\prime}}$) and $C_{0}\cup C_{b}\cup L_{z}$ (resp. $C_{0}^{\prime}\cup C^{\prime}_{b^{\prime}}\cup L_{z}$. Here are the corresponding results.
Theorem 1.7.
  1. (1)
  2. (2)
  3. (3)
Theorem 1.8.
  1. (1)
    The curves $C_{0}^{\prime}\cup L_{z}$ and $C_{1}^{\prime}\cup L_{z}$ are nearly free with exponents $(5,6)$ and $\tau(C_{0}^{\prime}\cup L_{z})=\tau(C_{1}^{\prime}\cup L_{z})=74$.
  2. (2)
    The curve $C^{\prime}_{b^{\prime}}\cup L_{z}$ is free with exponents $(5,5)$ and $\tau(C^{\prime}_{b^{\prime}}\cup L_{z})=75$.
  3. (3)
    The curve $C_{0}^{\prime}\cup C^{\prime}_{b^{\prime}}$ (resp. $C_{0}^{\prime}\cup C^{\prime}_{b^{\prime}}\cup L_{z}$) is nearly free with exponents $(9,11)$ and $\tau(C_{0}^{\prime}\cup C^{\prime}_{b^{\prime}})=270$, (resp. exponents $(9,12)$ and $\tau(C_{0}^{\prime}\cup C^{\prime}_{b^{\prime}}\cup L_{z})=300$).
In the pencils of curves ${\alpha}C_{0}+{\beta}C_{b}$ (resp. ${\alpha}C_{0}^{\prime}+{\beta}C_{b^{\prime}}^{\prime}$) there are 3 exceptional members (i.e., members more singular than the general member), namely $C_{0},C_{b}$ and $10L_{z}$ (resp. $C_{0}^{\prime}$, $C_{b^{\prime}}^{\prime}$. and $10L_{z}$). Therefore the freeness of the curve $C_{0}\cup C_{b}\cup L_{z}$ (resp. $C_{0}^{\prime}\cup C^{\prime}_{b^{\prime}}\cup L_{z}$) can be related to the results in [7], see also [9, 26].
Note that the lines $L_{x}:x=0$ and $L_{y}:y=0$ satisfy
Theorem 1.9.
Question 1.10.
It seems that this question is related to the very challenging question of finding numerical restrictions for Briançon-type polynomials, involving the number of special fibers and their geometry.
2. A new characterization of $MPOG$-curves
@@
Recall that the total Tjurina number $\tau(C)$ of a reduced plane curve $C:f=0$ is just the degree of its Jacobian ideal $J_{f}$, or equivalently the sum of the Tjurina numbers of all the singularities of $C$. The following result, due to du Plessis and Wall, see [17, Theorem 3.2] as well as [19] for an alternative approach, gives bounds on this Tjurina number $\tau(C)$.
@@
In the paper [8], the first author has given an alternative proof of Corollary 2.2 and has shown that a similar property holds for nearly free curves.
@@
2.4. The proof of Theorem 1.2
@@
If we assume that $C$ is an $MPOG$-curve, then the equality for $\tau(C)$ follows from [15, Proposition 2.1, (4)].
@@
a contradiction. It follows that $r\leq d/2$. Now [10, Theorem 1.2] implies that
@@
when $r<d/2$. In the case $r=d/2$, we have the following equalities, as in the discussion after [10, Theorem 1.2].
@@
since $\tau(C)=\tau(d,r)_{max}-2=3r^{3}-3r-1.$ Therefore the assumption $\tau(C)=\tau(d,r)_{max}-2$ implies in all cases that $\nu(C)=2$. Now we apply [15, Theorem 3.11] and conclude that $C$ is an $MPOG$-curve.
@@
does not determine uniquely the exponents of $C$. Indeed, such a curve may be a plus-one generated curve with exponents $(r,d-r,d-r+2)$, a 4-syzygy curve with exponents $(r,r,r,r)$ and degree $d=2r-1$ as in [15, Theorem 3.11], or maybe even other types of curves.
@@
The free and nearly free curves have strong relations with the rational cuspidal plane curves, see [13, 14]. The $MPOG$-curves are related with the rational nearly cuspidal curves, i.e. rational curves whose singularities are all irreducible except one singularity, which has two branches, see [15, Theorem 1.3, Conjecture 1.4, Corollary 5.8]. See also [21] for an alternative view on nearly free curves.

### v4

# Plus-one generated curves, Briançon-type polynomials and eigenscheme ideals

@@ Abstract.
We define the minimal plus-one generated curves and prove a result explaining why they are the closest relatives of the free curves, after the nearly free curves. Then we look at the projective closures of the general and of the special fibers of some Briançon-type polynomials constructed by E. Artal Bartolo, Pi. Cassou-Noguès and I. Luengo Velasco. They yield new examples of free, nearly free or minimal plus-one generated curves, as well as counter-examples to the conjecture saying that a supersolvable curve is free. In the final section we give a characterization of plus-one generated curves in terms of eigenscheme ideals, similar to the characterization of free curves given by R. Di Gennaro, G. Ilardi, R.M. Miró-Roig, H. Schenck and J. Vallès in a recent paper. Then we apply this result to the construction of minimal plus-one generated curves obtained by putting together at least two members in a pencil of curves related to Briançon-type polynomials.
@@ 1. Introduction
Let $S=\mathbb{C}[x,y,z]$ be the polynomial ring in three variables $x,y,z$ with complex coefficients, and let $C:f=0$ be a reduced curve of degree $d\geq 3$ in the complex projective plane $\mathbb{P}^{2}$. We denote by $J_{f}$ the Jacobian ideal of $f$, i.e., the homogeneous ideal in $S$ spanned by the partial derivatives $f_{x},f_{y},f_{z}$ of $f$, and by $M(f)=S/J_{f}$ the corresponding graded quotient ring, called the Jacobian (or Milnor) algebra of $f$. The $S$-module of derivations $Der(S)$ of the polynomial ring $S$ is a free module with a basis given by $\partial_{x}$, $\partial_{y}$ and $\partial_{z}$, that is any derivation $\theta\in Der(S)$ can be uniquely written as
(1.1) $\theta=a\partial_{x}+b\partial_{y}+c\partial_{z}$
with $a,b,c\in S$. When $a,b,c\in S_{m}$, we say that $\theta$ has degree $m$ and write $\deg\theta=m$. Consider the graded $S$-module of Jacobian syzygies of $f$ or, equivalently, the module of derivations killing $f$, namely
@@
The graded $S$-module
$D(f)=\{\theta\in Der(S)\ :\ \theta(f)\in(f)\}$
of derivations preserving $f$ has a direct sum decomposition in each degree $s$ given by
(1.3) $D(f)_{s}=S_{s-1}\cdot E\oplus D_{0}(f)_{s},$
for any integer $s$, where $E=x\partial_{x}+y\partial_{y}+z\partial_{z}$ is the Euler derivation and $D_{0}(f)$ is the submodule of derivations killing $f$ as in (1.2). We say that $C:f=0$ is an $m$-syzygy curve if the module $D_{0}(f)$ is minimally generated by $m$ homogeneous syzygies, say $\rho_{1},\rho_{2},\ldots,\rho_{m}$, of degrees $d_{j}=\deg\rho_{j}$ ordered such that
@@
The curve $C$ is free when $m=2$, since then $D_{0}(f)$ is a free module of rank 2, see for instance [26, 27, 28, 29]. A free curve $C$ is also characterized by the condition $d_{1}+d_{2}=d-1$, and hence in this case $r=d_{1}<d/2$, see [28]. A nearly free curve, as introduced in [15], is a 3-syzygy curve $C$ with exponents $(d_{1},d_{2},d_{3})$ such that $d_{1}+d_{2}=d$ and $d_{2}=d_{3}$. To simplify the notation, we say that $(d_{1},d_{2})$ are the exponents for the nearly free curve $C$. Finally, a plus-one generated curve is a 3-syzygy curve $C$ with exponents $(d_{1},d_{2},d_{3})$ such that $d_{1}+d_{2}=d$, see [1, 17]. We define the $\delta$-level of a plus-one generated curve by the formula
@@
These curves can be regarded as the closest neighbors of the free curves beyond the nearly free curves. Before stating the result, we recall some notions and results. The total Tjurina number $\tau(C)$ of a reduced plane curve $C:f=0$ is just the degree of its Jacobian ideal $J_{f}$, or equivalently the sum of the Tjurina numbers of all the singularities of $C$. The following result, due to du Plessis and Wall, see [20, Theorem 3.2] as well as [22] for an alternative approach, gives bounds on this Tjurina number $\tau(C)$.
@@
In the paper [10], the first author has given an alternative proof of Corollary 1.3 and has shown that the following similar property holds for nearly free curves.
@@
Our first main result in this paper is the following.
Theorem 1.5.
$\tau(C)=\tau(d,r)_{max}-2.$
$N(f)=I_{f}/J_{f}.$
$\nu(C)=\max_{j}\{n(f)_{j}\}$
as in [2].
Remark 1.6.
Note that $C$ is free if and only if $N(f)=0$, see for instance [28], and hence in this case $\nu(C)=0$, and $C$ is nearly free if and only if $\nu(C)=1$, see [15]. To prove Theorem 1.5, we show in fact that $\tau(C)=\tau(d,r)_{max}-2$ implies $\nu(C)=2$ and then use the classification of such curves given in [17, Theorem 3.11]. This result says that a curve $C$ with $\nu(C)=2$ is either an MPOG-curve, or a 4-syzygy curve with exponents $(r,r,r,r)$ and degree $d=2r-1$, which is in fact a maximal Tjurina curve of type $(d,r)=(2r-1,r)$ as defined above.
Let $s=xy+1$, $p=xs+1$ and $u=s^{2}+y$. Using this notation, we define
The polynomial $g$ up-to a change of coordinates $(x,y)\mapsto(-x,-y)$ was considered in an unpublished manuscript by J. Briançon, see [4, 5]. We recall that for a polynomial mapping $h:\mathbb{C}^{2}\to\mathbb{C}$ the set of atypical values is the minimal set $A(h)\subset\mathbb{C}$ such that $h$ induces a locally trivial topological fibration over the set $\mathbb{C}\setminus(A(h)\cup C(h))$, where $C(h)$ denotes the set of critical values of $h$. The atypical values for a function $h:\mathbb{C}^{2}\to\mathbb{C}$ come from jumping of the Milnor numbers of the singularities of $h$ at infinity, see [7, 21, 23]. The polynomials $g$ and $g^{\prime}$ have the following very interesting properties, see [4].
Theorem 1.7.
  1. (1)
  2. (2)
  3. (3)
    The set of atypical values for $g$ is $A(g)=\{0,b\}$, where $b=-\frac{16}{9}.$
  4. (4)
  5. (5)
When $h:\mathbb{C}^{2}\to\mathbb{C}$ is obtained from a linear form in $R$ by composition with a polynomial automorphism of $\mathbb{C}^{2}$, the freeness properties of the projective closure $D_{t}$ of any fiber $h^{-1}(t)$ are discussed in [13, Theorem 1.5]. It is natural to ask what happens when we replace such a polynomial $h$ by a Briançon-type polynomial.
For the polynomial $g$, we consider three associated plane curves $C_{t}$ in $\mathbb{P}^{2}$, which are the projective closure of the affine fibers $F_{t}=g^{-1}(t)$ for $t=0$, $t=b$ and $t=1$. The first two fibers correspond to the atypical values in $A(g)$, and the third one is a general fiber of $g$. For the convenience of the reader, the defining equations for these curves, obtained by homogenization starting with the first equation in (1.5), are given in (3.1) and (3.2). Similarly, for the polynomial $g^{\prime}$, we consider three associated plane curves $C_{t}^{\prime}$ in $\mathbb{P}^{2}$, which are the projective closure of the affine fibers $g^{\prime-1}(t)$ for $t=0$, $t=b^{\prime}$ and $t=1$. The corresponding equations are given in (3.3) and (3.4).
Our results about the freeness properties of these projective plane curves are given in Proposition 3.3 and Proposition 3.4.
Since the line at infinity $L_{z}:z=0$ intersects the curves $C_{t}$ only in two points, the general principle explained in [14] tells us that the curves $C_{t}\cup L_{z}$ (resp. $C_{t}^{\prime}\cup L_{z}$) of degree $11$ may also enjoy (sometimes better) freeness properties. See also Step 3 in the proof of Theorem 1.5 in [13] for such a situation. We consider also the curve arrangements $C_{0}\cup C_{b}$ (resp. $C_{0}^{\prime}\cup C^{\prime}_{b^{\prime}}$) and $C_{0}\cup C_{b}\cup L_{z}$ (resp. $C_{0}^{\prime}\cup C^{\prime}_{b^{\prime}}\cup L_{z}$. The corresponding results are given in Propositions 3.6, 3.7 and 3.8.
At the end of Section 3, we construct two counter-examples involving the curve $C_{0}$ to the conjecture [14, Conjecture 1.10], which asks whether any supersolvable curve is free. In the first example the modular point is on $C_{0}$, while in the second example the modular point is not on $C_{0}$, see Remark 3.11, where we recall the definitions of all the notions involved here.
Question 1.8.
It seems that this question is related to the very challenging question of finding numerical restrictions for Briançon-type polynomials, involving their geometry.
Finally we give a characterization of plus-one generated curves in terms of eigenscheme ideals, similar to the following characterization of free curves given by R. Di Gennaro, G. Ilardi, R.M. Miró-Roig, H. Schenck and J. Vallès in a recent paper [9]. For a derivation $\theta$ as in (1.1), we consider the ideal $I_{\theta}\subset S$ generated by the three 2-minors of the matrix having as a first row $(x,y,z)$ and as a second row the coefficients $(a,b,c)$ of $\theta$. The eigenscheme $\Gamma_{\theta}$ associated to $\theta$ is the subscheme in $\mathbb{P}^{2}$ defined by the ideal $I_{\theta}$. When the eigenscheme $\Gamma_{\theta}$ is 0-dimensional, then the ideal $I_{\theta}$ is saturated, see [6, Proposition 3.1]. The following result was obtained in [9], see Theorem 2.5.
Theorem 1.9.
Let $C:f=0$ be a reduced curve in $\mathbb{P}^{2}$ of degree $d$. Let $\theta\in D_{0}(f)$ be a non trivial derivation of degree $r$. We assume that $12r\leq d-$ 1 and that the eigenscheme $\Gamma_{\theta}$ is 0-dimensional. Then $C$ is a free curve with exponents $(r,d-r-1)$ if and only if $f\in I_{\theta}$.
The condition $12r\leq d-$ 1 is added here since we like our exponents $(d_{1},d_{2})$ to verify the property $d_{1}\leq d_{2}$. See also Theorem 1.10 $(1)$ below. In view of the direct sum decomposition, one may even take $\theta\in D(f)_{r}$, as in [9, Theorem 2.5], and the same result holds. A key remark is that $f\in I_{\theta}$ if and only if the ideal quotient
$K_{\theta,f}=I_{\theta}:(f)=\{h\in S\ :\ hf\in I_{\theta}\}$
is the whole ring $S$, where $(f)$ denotes the principal ideal in $S$ generated by $f$. From this new point of view, Theorem 1.9 can be modified to cover the case of plus-one generated curves, as our second main result shows.
Theorem 1.10.
Let $C:f=0$ be a reduced curve in $\mathbb{P}^{2}$ of degree $d$. Let $\theta\in D_{0}(f)$ be a non trivial derivation of degree $r$. We assume that $12r\leq d-$ 1 and that the eigenscheme $\Gamma_{\theta}$ is 0-dimensional. Then the following hold.
  1. (1)
    $r={\rm mdr}(f)$.
  2. (2)
    $C$ is a plus-one generated curve if and only if the ideal quotient $K_{\theta,f}=I_{\theta}:(f)$ is a proper ideal having the degree one component $(K_{\theta,f})_{1}$ nonzero.
If the properties in $(2)$ hold, then the ideal $K_{\theta,f}$ has two generators, a linear form $\ell\in S_{1}$ and a homogeneous polynomial $h\in S_{e}$, with $e\geq 1$, and the exponents of the curve $C$ are $(r,d-r,d-r-1+e)$. In particular, $C$ is nearly free if and only if $e=1$ and $C$ is an MPOG-curve if and only if $e=2$.
This result is applied to the construction of minimal plus-one generated curves obtained by putting together some members in a pencil of curves, see Corollary 5.5 for the general situation, and Proposition 5.6, for an explicit construction related to Briançon-type polynomials. To our knowledge, this is the first example where putting together the most singular members in a pencil of plane curves does not produce a free curve, see also [9, Remark 3.1].
The first author thanks Pierrette Cassou-Noguès for useful discussion on this paper, reflected in part in Remark 3.10.
2. The proof of Theorem 1.5
@@
If we assume that $C$ is an $MPOG$-curve, then the equality for $\tau(C)$ follows from [17, Proposition 2.1, (4)].
@@
a contradiction. It follows that $r\leq d/2$. Now [12, Theorem 1.2] implies that
@@
when $r<d/2$. In the case $r=d/2$, we have the following equalities, as in the discussion after [12, Theorem 1.2].
@@
since $\tau(C)=\tau(d,r)_{max}-2=3r^{3}-3r-1.$ Therefore the assumption $\tau(C)=\tau(d,r)_{max}-2$ implies in all cases that $\nu(C)=2$. Now we apply [17, Theorem 3.11] and conclude that $C$ is an $MPOG$-curve.
@@
does not determine uniquely the exponents of $C$. Indeed, such a curve may be a plus-one generated curve with exponents $(r,d-r,d-r+2)$, a 4-syzygy curve with exponents $(r,r,r,r)$ and degree $d=2r-1$ as in [17, Theorem 3.11], or maybe even other types of curves.
@@
The free and nearly free curves have strong relations with the rational cuspidal plane curves, see [15, 16]. The $MPOG$-curves are related with the rational nearly cuspidal curves, i.e. rational curves whose singularities are all irreducible except one singularity, which has two branches, see [17, Theorem 1.3, Conjecture 1.4, Corollary 5.8]. See also [25] for an alternative view on nearly free curves.